OpenStudy (anonymous):

A car dealer received a shipment of cars, half of which were black, with the remainder consisting of equal numbers of blue, silver, and white cars. During the next month, 70 percent of the black cars, 80 percent of the blue cars, 30 percent of the silver cars, and 40 percent of the white cars were sold. What percent of the cars in the shipment were sold during that month.

OpenStudy (anonymous):

i dont need the answer i need to know how to do it....

OpenStudy (kropot72):

 You need to calculate the number of cars sold for each color. Then find the total number sold. The number of black cars sold is \[\large \frac{n}{2}\times\frac{0.7}{1}=0.35n\]

OpenStudy (kropot72):



OpenStudy (kropot72):

The total number of cars in the shipment = n